OpenStudy (anonymous):

Will give medal! Decide which part of the quadratic formula tells you whether the quadratic equation can be solved by factoring. −b b^2 − 4ac 2a Use the part of the quadratic formula that you chose above and find its value, given the following quadratic equation: 4x^2 + 6x + 2 = 0

OpenStudy (anonymous):

(x+1) (4x+2) ?

OpenStudy (anonymous):

(b^2 - 4ac) is called the discriminant. If discriminant is positive and a perfect square then the quadratic equation can be solved by factoring.

OpenStudy (anonymous):

so did 1 factor correctly?

OpenStudy (anonymous):

woh let's not get ahead of ourselves

OpenStudy (anonymous):

4x2 + 6x + 2 = 0 First let us find the discriminant: b^2 - 4ac 6^2 - (4)(4)(2) = 36 - 32 = 4 Discriminant is positive and a perfect square and therefore the quadratic equation can be solved by factoring. Here I don't think they are asking you to factor it. They are saying: "Use the part of the quadratic formula that you chose above and find its value" The part you chose for the previous question was: b^2 - 4ac and its value is 4. But if you need to factor it then: ...

OpenStudy (anonymous):

4x^2 + 6x + 2 = 0 4x^2 + 4x + 2x + 2 = 0 4x(x + 1) + 2(x + 1) = 0 (4x + 2)(x + 1) = 0 we can further factor 2 out of the first factor: 2(2x + 1)(x + 1) = 0

OpenStudy (anonymous):

so the answer would only be the number that represents that certain part of the quadratic equation?

OpenStudy (anonymous):

yes -1/2 & -1

OpenStudy (anonymous):

so the answer to the question is 4?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

but it asks for the part of the quadratic equation that 1 chose?

OpenStudy (anonymous):

oh yes sorry

OpenStudy (anonymous):

yep 4

OpenStudy (anonymous):

okay thanks